One of the most recent transuranium elements oganesson-294 (Og - 294), atomic number 1 Part A Complete the following equation

one of the most recent transuranium elements oganesson-294(Og-294), atomic number 118, was prepared by bombarding califomium-249 with another isotope
Answer #1

let ?? = x 24.9 294 ?? & Cf Og + 3 n 98 118 Balance atomie mass A + 249 = 294 + 3(1) A + 249= 294 + 3 A = 297- 249 A = 48 num